# Hyperspectral Pansharpening: Critical Review, Tools and Future Perspectives

### v1

@@ III-B3 Regression-based QNR
The $D_{\lambda}$ and $D_{S}$ indexes measure two different aspects of the quality of pansharpened images. An algorithm that minimizes one index may cause the other to overgrow, leading to poor overall quality. Therefore, they should be taken into account jointly, looking for the best trade-off between spatial and spectral distortion. This is the objective of the Regression-based QNR (RQNR) index defined as follows (see [89] for additional deteils):
@@ IV Benchmarking datasets
TABLE IV: Review of HS pansharpening datasets. Resolutions and sizes refer to the target fusion image.
Work Instrument - Scene HS PAN $R$ Bands Spectral Resolution Size Train/Valid. Test
(used/all) range patches image
HyperPNN [34] HYDICE - Washington DC Mall $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 5 191/210 0.4–2.4 $\mu$m 2 m 1208$\times$307 n.a.$\times$(11$\times$11)(o) 256$\times$128
@@
AVIRIS - Salinas Valley $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 5 204/224 0.4–2.5 $\mu$m 3.7 m 510$\times$215 n.a.$\times$(11$\times$11)(o) 256$\times$128
HSpeNet [38] HYDICE - Washington DC Mall $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 5 191/210 0.4–2.4 $\mu$m 2 m 1208$\times$307 4284$\times$(11$\times$11)(o) 200$\times$200
AVIRIS - Moffett Field $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 5 176/224 0.4–2.5 $\mu$m 20 m 395$\times$185 1645$\times$(11$\times$11)(o) 150$\times$150
ROSIS - Pavia University $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 5 103/115 0.4–0.9 $\mu$m 1.3 m 610$\times$340 3120$\times$(11$\times$11)(o) 200$\times$200
EO-1/ALI - Halls Creek $\mathbf{H}_{\downarrow}$ $\mathbf{P}_{\downarrow}$ 3 171/242 0.4–2.5 $\mu$m 30 m 1161$\times$189 n.a.$\times$(11$\times$11)(o) 120$\times$120
EO-1/ALI - Halls Creek real real 3 171/242 0.4–2.5 $\mu$m 10 m 3483$\times$567 n.a.$\times$(11$\times$11)(o) n.a.
DHP-DARN [35] CAVE (real world scenes) $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 4 31/31 0.4–0.7 $\mu$m n.a. 32$\times$(512$\times$512) 5632$\times$(32$\times$32) 10$\times$(512$\times$512)
ROSIS - Pavia Center $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 4 102/115 0.4–0.9 $\mu$m 1.3 m 1096$\times$715 425$\times$(32$\times$32) 7$\times$(160$\times$160)
EO-1 - Botswana $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 3 145/242 0.4–2.5 $\mu$m 30 m 1496$\times$256 126$\times$(32$\times$32) 6$\times$(120$\times$120)
EO-1/ALI - Los Angeles real real 3 145/242 0.4–2.5 $\mu$m 10 m 360$\times$360 360$\times$360
DIP-HyperKite [36] ROSIS - Pavia Center $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 4 102/115 0.4–0.9 $\mu$m 1.3 m 1096$\times$715 17$\times$(160$\times$160) 7$\times$(160$\times$160)
MV.C VNIR - Chikusei $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 4 128/512 0.4–1.0 $\mu$m 2.5 m 2517$\times$2335 61$\times$(256$\times$256) 20$\times$(256$\times$256)
EO-1 - Botswana $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 3 145/242 0.4–2.5 $\mu$m 30 m 1496$\times$256 14$\times$(120$\times$120) 6$\times$(120$\times$120)
EO-1/ALI - Los Angeles real real 3 145/242 0.4–2.5 $\mu$m 10 m 360$\times$360 360$\times$360
HyperDSNet [62] HYDICE - Washington DC Mall $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 4 191/210 0.4–2.4 $\mu$m 2 m 1208$\times$307 1024$\times$(64$\times$64)(o) 4$\times$(128$\times$128)
ROSIS - Pavia Center $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 4 102/115 0.4–0.9 $\mu$m 1.3 m 1096$\times$715 1680$\times$(64$\times$64)(o) 2$\times$(400$\times$400)
EO-1 - Botswana $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 4 145/242 0.4–2.5 $\mu$m 30 m 1496$\times$256 967$\times$(64$\times$64)(o) 4$\times$(128$\times$128)
PRISMA - Bologna $\mathbf{H}_{\downarrow}$ $\mathbf{P}_{\downarrow}$ 6 69/239 0.4–2.5 $\mu$m 30 m 400$\times$400 816$\times$(60$\times$60)(o)
PRISMA - Bologna real real 6 69/239 0.4–2.5 $\mu$m 5 m 2400$\times$2400 2$\times$(240$\times$240)
R-PNN [37] PRISMA - Barcelona $\mathbf{H}_{\downarrow}$ $\mathbf{P}_{\downarrow}$ 6 59/239 0.4–2.5 $\mu$m 30 m 900$\times$900 900$\times$900
PRISMA - Milan $\mathbf{H}_{\downarrow}$ $\mathbf{P}_{\downarrow}$ 6 73/239 0.4–2.5 $\mu$m 30 m 900$\times$900 900$\times$900
PRISMA - Prato real real 6 73/239 0.4–2.5 $\mu$m 5 m 2400$\times$2400 100$\times$(240$\times$240)(f)
PRISMA - Bologna real real 6 69/239 0.4–2.5 $\mu$m 5 m 2400$\times$2400 2400$\times$2400
PRISMA - Florence real real 6 63/239 0.4–2.5 $\mu$m 5 m 2400$\times$2400 2400$\times$2400
PCA-Z-PNN [50] PRISMA - Milan $\mathbf{H}_{\downarrow}$ $\mathbf{P}_{\downarrow}$ 6 73/239 0.4–2.5 $\mu$m 30 m 900$\times$900 900$\times$900(v)
PRISMA - Barcelona $\mathbf{H}_{\downarrow}$ $\mathbf{P}_{\downarrow}$ 6 59/239 0.4–2.5 $\mu$m 30 m 900$\times$900 900$\times$900
ROSIS - Pavia University $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 6 103/115 0.4–0.9 $\mu$m 1.3 m 610$\times$340 606$\times$336
PRISMA - Prato real real 6 73/239 0.4–2.5 $\mu$m 5 m 2400$\times$2400 2400$\times$2400(v)
PRISMA - Bologna real real 6 69/239 0.4–2.5 $\mu$m 5 m 2400$\times$2400 2400$\times$2400
PRISMA - Florence real real 6 63/239 0.4–2.5 $\mu$m 5 m 2400$\times$2400 2400$\times$2400
@@
Before presenting the proposed benchmarking dataset, it is worth reviewing the most popular datasets used for HS pansharpening evaluation. Since this topic is particularly important for DL-based methods, we limit attention to the latter, and in particular to those considered in the toolbox. In any case, these datasets are quite representative of the overall literature (DL or not) on HS pansharpening. They are gathered in Tab. IV, with all the details on their characteristics and use.
@@
For each dataset, the HS and PAN columns indicate whether these two components are real (as distributed by the provider) or synthetic. The latter are obtained through spatial resolution downgrading ($\mathbf{H}_{\downarrow}$, $\mathbf{P}_{\downarrow}$) or by averaging the HS bands, limiting the scope to the visible spectrum ($\langle\mathbf{H}\rangle$). In most cases, the use of synthetic components is unavoidable because the original image is not multi-resolution and only the HS component is given, with no PAN. Sometimes, instead, truly multiresolution data are available, like the EO-1/ALI combination or the PRISMA images. In these cases, two options are possible: i) downgrade the resolution of both HS and PAN anyway, in order to use supervised training and full-reference quality metrics; ii) keep the original data to preserve information, and use some forms of unsupervised training together with no-reference quality metrics.
@@
Continuing to scan the columns of Tab. IV, we first find the resolution ratio, $R$, which goes from 3 to 6, the most challenging case. Actually, this value is arbitrary for all datasets without a real PAN component. Then, the number of bands (used/all) points out the critical issue of image structure variability that occurs not only for different sensors but also for different scenes acquired with the same instrument. The spectral range covered is rather uniform, going from the visible (0.4 $\mu$m on) to the short-wave infrared (up to 2.4-2.5 $\mu$m), with the exception of ROSIS and MV.C VNIR, which cover up to the near-infrared and CAVE, limited to the visible spectrum. The physical size of the pixels, known as spatial resolution or ground sample distance (GSD), is also quite variable, ranging from 1.3 m up to 30 m, referring to the pansharpened image domain, and accounting also for resolution downgrade if any. Sometimes, the resolution is very high. This is due to the use of airborne instruments (e.g., AVIRIS) operated at a much lower altitude than satellite ones. This is a further source of variability that must be taken into account.
@@
The analysis of Tab. IV shows a rather puzzling scenario. Different groups use different test images, so the results are not immediately comparable with the prior art. Furthermore, they train models on images characterized by a number of bands that changes from case to case, so most of the trained models cannot be used on new images with a different number of bands. This, in turn, prevents the generalization ability of the methods to be verified. Finally, training sets are often just too small to ensure correct training.
@@ IV-A The proposed dataset
A major contribution of this work is to provide a dataset that meets all the above requirements. Considering the restrictive data-sharing policies widespread in the remote sensing field, we decided to use the PRISMA (PRecursore IperSpettrale della Missione Applicativa) images, shared on-demand by the Italian Space Agency (ASI) for research purposes only. In particular, a set of 16 images has been carefully selected and organized as summarized in Tab. V, including both the full-resolution and reduced-resolution versions. The first set of 12 images is reserved for training and validation purposes (part A of the table). All images are rather large, 3600$\times$3600 pixels at the target 5m resolution, which allows us to extract 10 or more large tiles from each of them, with size 1152$\times$1152 at full resolution and 192$\times$192 after the 6$\times$ decimation needed for RR training. As always happens with HS images, not all bands have sufficient quality for further processing. Therefore, out of the 239 bands available, we identified a subset of 159 bands that have good quality in all training and test images. A second set of 4 images, is kept for testing (part B) and form the strict-sense benchmarking dataset. Again, both full and reduced-resolution versions are given. In the former case, only 1200$\times$1200 crops are used, to limit complexity, while the complete 600$\times$600 RR images are used in the latter case. The images have been acquired over a period of more than 3 years, all over the world, and display a wide variety of land covers. In Fig. 2 we show an RGB composition of the original HS images. For RR tests these are used in their full size, whereas smaller tiles (red solid line-boxes) are considered for FR tests to limit the computation time. In fact, in the FR space each HS test image would get a size of 3600$\times$3600$\times$159, once interpolated to match the target spatial size, giving rise to a considerable volume. The tiles are instead 200$\times$200 pixels wide, hence 1200$\times$1200 with interpolation, i.e. 4 times larger images compared to the RR test images. Red and green dashed line-boxes highlight the crops sampled for visualization purposes in the experimental section for FR and RR tests, respectively.
@@
TABLE V: PRISMA benchmark datasets: (A) Training, (B) Test.

### v2

@@ III-B3 Regression-based QNR
The $D_{\lambda}$ and $D_{S}$ indexes measure two different aspects of the quality of pansharpened images. An algorithm that minimizes one index may cause the other to overgrow, leading to poor overall quality. Therefore, they should be taken into account jointly, looking for the best trade-off between spatial and spectral distortion. This is the objective of the Regression-based QNR (RQNR) index defined as follows (see [98] for additional details):
@@
It is worth to notice the similarity between RQNR and the Hybrid Quality with No Reference (HQNR) index [111] commonly used in the MS case which differs from the former for a different formulation of the spatial distortion term $D_{S}$. The spatial distortion index in HQNR treats any alterations in the relationship between MS and PAN across different resolution scales as spatial distortions. This approach can lead to coupling effects between the two distortion metrics. In cases, as it is the case here, where many bands lack spectral overlap with the PAN image, spectral distortion becomes dominant and can impact the measurement of spatial distortion in HQNR. In contrast, the spatial metric employed in RQNR mitigates these coupling effects and is therefore preferred [98].
@@ IV Benchmarking datasets
TABLE IV: Popular datasets for HS pansharpening.
Dataset Name Bands Spectral Size Paper Repository
(used/all) Range $N\times\left(H\times W\right)$
HYDICE - Washington DC Mall 191/210 0.4–2.4 $\mu$m 1$\times$(1208$\times$307) a
MV.C VNIR - Chikusei 128/512 0.4–1.0 $\mu$m 1$\times$(2517$\times$2335) [112] b
ROSIS - Pavia Center 102/115 0.4–0.9 $\mu$m 1$\times$(1096$\times$715) c
ROSIS - Pavia University 103/115 0.4–0.9 $\mu$m 1$\times$(610$\times$340) c
AVIRIS - Moffet Field 176/224 0.4–2.5 $\mu$m 1$\times$(395$\times$185) d
AVIRIS - Salinas Valley 204/224 0.4–2.5 $\mu$m 1$\times$(510$\times$215) c
EO-1 - Botswana 145/242 0.4–2.5 $\mu$m 1$\times$(1496$\times$256) c
EO-1/ALI - Halls Creek 171/242 0.4–2.5 $\mu$m 1$\times$(3483$\times$567)
EO-1/ALI - Los Angeles 145/242 0.4–2.5 $\mu$m 1$\times$(360$\times$360)
CAVE 31/31 0.4–0.7 $\mu$m 32$\times$(512$\times$512) [113] e
PRISMA - Barcelona 59/239 0.4–2.5 $\mu$m 1$\times$(900$\times$900) [4] f
PRISMA - Milan 73/239 0.4–2.5 $\mu$m 1$\times$(900$\times$900) [4] f
PRISMA - Prato 73/239 0.4–2.5 $\mu$m 1$\times$(2400$\times$2400)
PRISMA - Bologna 69/239 0.4–2.5 $\mu$m 1$\times$(2400$\times$2400) [4] f
PRISMA - Florence 63/239 0.4–2.5 $\mu$m 1$\times$(2400$\times$2400) [4] f
@@
bhttps://naotoyokoya.com/Download.html
chttps://www.ehu.eus/ccwintco/index.php/Hyperspectral_Remote_Sensing_Scenes
dhttps://aviris.jpl.nasa.gov/data/free_data.html
ehttps://cave.cs.columbia.edu/projects/categories/project?cid=Computational+Imaging&pid=
Generalized+Assorted+Pixel+Camera
fhttps://openremotesensing.net/knowledgebase/panchromatic-and-hyperspectral-image-fusion
-outcome-of-the-2022-whispers-hyperspectral-pansharpening-challenge
TABLE V: Review of HS pansharpening datasets used by DL methods. Resolutions and sizes refer to the target fusion image.
Work Instrument - Scene HS PAN $R$ Resolution Train/Valid. Patches Test Patches
$N\times\left(H\times W\right)$ $N\times\left(H\times W\right)$
HyperPNN [34] HYDICE - Washington DC Mall $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 5 2 m n.a.$\times$(11$\times$11)(o) 1$\times$256$\times$128
@@
AVIRIS - Salinas Valley $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 5 3.7 m n.a.$\times$(11$\times$11)(o) 1$\times$(256$\times$128)
HSpeNet [38] HYDICE - Washington DC Mall $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 5 2 m 4284$\times$(11$\times$11)(o) 1$\times$(200$\times$200)
AVIRIS - Moffett Field $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 5 20 m 1645$\times$(11$\times$11)(o) 1$\times$(150$\times$150)
ROSIS - Pavia University $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 5 1.3 m 3120$\times$(11$\times$11)(o) 1$\times$(200$\times$200)
EO-1/ALI - Halls Creek $\mathbf{H}_{\downarrow}$ $\mathbf{P}_{\downarrow}$ 3 30 m n.a.$\times$(11$\times$11)(o) 1$\times$(120$\times$120)
EO-1/ALI - Halls Creek real real 3 10 m n.a.$\times$(11$\times$11)(o) n.a.
DHP-DARN [35] CAVE (real world scenes) $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 4 n.a. 5632$\times$(32$\times$32) 10$\times$(512$\times$512)
ROSIS - Pavia Center $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 4 1.3 m 425$\times$(32$\times$32) 7$\times$(160$\times$160)
EO-1 - Botswana $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 3 30 m 126$\times$(32$\times$32) 6$\times$(120$\times$120)
EO-1/ALI - Los Angeles real real 3 10 m 1$\times$360$\times$360
DIP-HyperKite [36] ROSIS - Pavia Center $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 4 1.3 m 17$\times$(160$\times$160) 7$\times$(160$\times$160)
MV.C VNIR - Chikusei $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 4 2.5 m 61$\times$(256$\times$256) 20$\times$(256$\times$256)
EO-1 - Botswana $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 3 30 m 14$\times$(120$\times$120) 6$\times$(120$\times$120)
EO-1/ALI - Los Angeles real real 3 10 m 1$\times$(360$\times$360)
HyperDSNet [62] HYDICE - Washington DC Mall $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 4 2 m 1024$\times$(64$\times$64)(o) 4$\times$(128$\times$128)
ROSIS - Pavia Center $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 4 1.3 m 1680$\times$(64$\times$64)(o) 2$\times$(400$\times$400)
EO-1 - Botswana $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 4 30 m 967$\times$(64$\times$64)(o) 4$\times$(128$\times$128)
PRISMA - Bologna $\mathbf{H}_{\downarrow}$ $\mathbf{P}_{\downarrow}$ 6 30 m 816$\times$(60$\times$60)(o)
PRISMA - Bologna real real 6 5 m 2$\times$(240$\times$240)
R-PNN [37] PRISMA - Barcelona $\mathbf{H}_{\downarrow}$ $\mathbf{P}_{\downarrow}$ 6 30 m 1$\times$(900$\times$900)
PRISMA - Milan $\mathbf{H}_{\downarrow}$ $\mathbf{P}_{\downarrow}$ 6 30 m 1$\times$(900$\times$900)
PRISMA - Prato real real 6 5 m 100$\times$(240$\times$240)(f)
PRISMA - Bologna real real 6 5 m 1$\times$(2400$\times$2400)
PRISMA - Florence real real 6 5 m 1$\times$(2400$\times$2400)
PCA-Z-PNN [50] PRISMA - Milan $\mathbf{H}_{\downarrow}$ $\mathbf{P}_{\downarrow}$ 6 30 m 1$\times$(900$\times$900)(v)
PRISMA - Barcelona $\mathbf{H}_{\downarrow}$ $\mathbf{P}_{\downarrow}$ 6 30 m 1$\times$(900$\times$900)
ROSIS - Pavia University $\mathbf{H}_{\downarrow}$ $\langle\mathbf{H}\rangle$ 6 1.3 m 1$\times$(606$\times$336)
[…]
PRISMA - Bologna real real 6 5 m 1$\times$(2400$\times$2400)
PRISMA - Florence real real 6 5 m 1$\times$(2400$\times$2400)
[…]
@@
Before presenting the proposed benchmarking dataset, it is worth reviewing the most popular datasets used for HS pansharpening evaluation. Since this topic is particularly important for DL-based methods, we limit attention to the latter, and in particular to those implemented for the toolbox. In any case, these datasets are quite representative of the overall literature (DL or not) on HS pansharpening. Tab. IV summarizes the main features of such datasets (bands, spectral range, size, documentation, availability), while Tab. V provides useful information about their use by different research teams. For each work/dataset, the HS and PAN columns indicate whether these two components are real (as distributed by the provider) or synthetic. The latter are obtained through spatial resolution downgrading ($\mathbf{H}_{\downarrow}$, $\mathbf{P}_{\downarrow}$) or by averaging the HS bands, limiting the scope to the visible spectrum ($\langle\mathbf{H}\rangle$). In most cases, the use of synthetic components is unavoidable because the original image is not multi-resolution and only the HS component is given, with no PAN. Sometimes, instead, truly multiresolution data are available, like the EO-1/ALI combination or the PRISMA images. In these cases, two options are possible: i) downgrade the resolution of both HS and PAN anyway, in order to use supervised training and full-reference quality metrics; ii) keep the original data to preserve information, and use some forms of unsupervised training together with no-reference quality metrics.
@@
Continuing to scan the columns of Tab. V, we first find the resolution ratio, $R$, which goes from 3 to 6, the most challenging case. Actually, this value is arbitrary for all datasets without a real PAN component. Then, the number of bands (used/all) points out the critical issue of image structure variability that occurs not only for different sensors but also for different scenes acquired with the same instrument. The spectral range covered is rather uniform, going from the visible (0.4 $\mu$m on) to the short-wave infrared (up to 2.4-2.5 $\mu$m), with the exception of ROSIS and MV.C VNIR, which cover up to the near-infrared and CAVE, limited to the visible spectrum. The physical size of the pixels, known as spatial resolution or ground sample distance (GSD), is also quite variable, ranging from 1.3 m up to 30 m, referring to the pansharpened image domain, and accounting also for resolution downgrade if any. Sometimes, the resolution is very high. This is due to the use of airborne instruments (e.g., AVIRIS) operated at a much lower altitude than satellite ones. This is a further source of variability that must be taken into account.
@@
The analysis of Tab. V shows a rather puzzling scenario. Different groups use different test images, so the results are not immediately comparable with the prior art. Furthermore, they train models on images characterized by a number of bands that changes from case to case, so most of the trained models cannot be used on new images with a different number of bands. This, in turn, prevents the generalization ability of the methods to be verified. Finally, training sets are often just too small to ensure correct training.
@@ IV-A The proposed dataset
A major contribution of this work is to provide a dataset that meets all the above requirements. Considering the restrictive data-sharing policies widespread in the remote sensing field, we decided to use the PRISMA (PRecursore IperSpettrale della Missione Applicativa) images, shared on-demand by the Italian Space Agency (ASI) for research purposes only. In particular, a set of 16 images has been carefully selected and organized as summarized in Tab. VI, including both the full-resolution and reduced-resolution versions. The first set of 12 images is reserved for training and validation purposes (part A of the table). All images are rather large, 3600$\times$3600 pixels at the target 5m resolution, which allows us to extract 10 or more large tiles from each of them, with size 1152$\times$1152 at full resolution and 192$\times$192 after the 6$\times$ decimation needed for RR training. As always happens with HS images, not all bands have sufficient quality for further processing. Therefore, out of the 239 bands available, we identified a subset of 159 bands that have good quality in all training and test images. A second set of 4 images, is kept for testing (part B) and form the strict-sense benchmarking dataset. Again, both full and reduced-resolution versions are given. In the former case, only 1200$\times$1200 crops are used, to limit complexity, while the complete 600$\times$600 RR images are used in the latter case. The images have been acquired over a period of more than 3 years, all over the world, and display a wide variety of land covers. In Fig. 2 we show an RGB composition of the original HS images. For RR tests these are used in their full size, whereas smaller tiles (red solid line-boxes) are considered for FR tests to limit the computation time. In fact, in the FR space each HS test image would get a size of 3600$\times$3600$\times$159, once interpolated to match the target spatial size, giving rise to a considerable volume. The tiles are instead 200$\times$200 pixels wide, hence 1200$\times$1200 with interpolation, i.e. 4 times larger images compared to the RR test images. Red and green dashed line-boxes highlight the crops sampled for visualization purposes in the experimental section for FR and RR tests, respectively.
@@
TABLE VI: PRISMA benchmark datasets: (A) Training, (B) Test.